OpenStudy (anonymous):

A solution of 78% alcohol is to be mixed with a solution of 23% alcohol to form 165 L of a 58% solution. How many liters of the 78% solution must be used?

OpenStudy (anonymous):

lets put the number of L of 78% alcohol as \(x\) so the number of L of 23% solution must be \(165-x\) since the total L is 165

OpenStudy (anonymous):

then you will have \(.78x+.23(165-x)\) liters of alcohol, which you want to be equal to \(.58\times 165=95.7\)

OpenStudy (anonymous):

so set \[.78x+.23(165-x)=95.7\] and solve for \(x\) although it is easier to multiply both sides by 100 and solve \[78x+23(165-x)=9570\]

OpenStudy (anonymous):

you just confuse me dude

OpenStudy (anonymous):

all steps are there and explained in the best english i know how. since you will probably have to do more of these, it might be a good idea to follow the steps and see if you understand them. perhaps there is a better (clearer) explanation in your text

OpenStudy (anonymous):

105 is what i got?

OpenStudy (anonymous):

or is it 57.178?

OpenStudy (anonymous):

help me please!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!